The LCM of numbers 15,70,105 is:can you give this anwer?
Answer by Edwin McCravy(20086)

15,70,105
Break them down into primes, like this (you can use factor trees).

 15 = 3*5,  70 = 2*5*7,  105 = 3*5*7

Now list like factors like this, so that like factor line up
vertically in columns:

 15 =   3*5
 70 = 2  *5*7
105 =   3*5*7

Then draw a line underneath:

 15 =   3*5
 70 = 2  *5*7
105 =   3*5*7
-------------

Now under that line bring down the factor of each column:

 15 =   3*5
 70 = 2  *5*7
105 =   3*5*7
-------------
      2*3*5*7

And multiply those together and get 210.  That's the LCM

Edwin